# Passage-traversing optimal path planning with sampling-based algorithms

### v1

@@ 4.3 Constrained passage width PTOPP
The last category of PTOPP focuses on problems with constraints on passed passage widths, termed constrained passage width PTOPP (CPW-PTOPP). The most representative case is that there exists a minimum admissible passage width $\underline{\varepsilon}_{p}$ to accommodate robots, similar to the preset clearance requirement. Technically, this is solvable by labeling passage regions of widths smaller than $\varepsilon_{p}$ as obstacles to prohibit traversing. In PTOPP framework, such constraints are easily imposed on path costs without the labeling processing above. Firstly, CPW-PTOPP can be easily integrated into shortest path planning. If the path passes a passage narrower than $\underline{\varepsilon}_{p}$, a penalty is imposed by adding a large value $K_{p}$ to the path cost. The path cost is expressed as
@@
where $|\mathbf{p}_{\sigma}|_{\underline{\varepsilon}_{p}}$ is the number of elements in $\mathbf{p}_{\sigma}$ smaller than $\underline{\varepsilon}_{p}$. The cost compatibility is easy to verify.
Next, constrained passage width could be enforced in PTOPP categories by truncating passage widths smaller than $\underline{\varepsilon}_{p}$ to be a negative value (e.g., $-1/\|l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})\|$ here) so that passage widths in $\mathbf{p}_{\sigma}$ are updated as follows
@@
This causes the path cost $f_{c}(\sigma)$ to soar after passing passages narrower than $\underline{\varepsilon}_{p}$, excluding them from the optimal path’s route. $f_{c}(\sigma)$ refers to any cost penalizing small passage widths. Therefore, CPW-PTOPP is adapted from previous PTOPP categories by simply applying (26) to $\mathbf{p}_{\sigma}$, making constraints on passage widths quite convenient to impose. Note that both MPW-PTOPP and GPW-PTOPP maximize the smallest passage width $f_{p}(\sigma)$ under the proposed path costs. If $f_{p}(\sigma)>\underline{\varepsilon}_{p}$, (26) does not change the optimal path. If $f_{p}(\sigma)\leq\underline{\varepsilon}_{p}$, (26) is beneficial for it helps select wider passages and decrease the number of passages narrower than $\underline{\varepsilon}_{p}$ in order to reach a smaller cost.
@@ 5 Primitive procedures in PTOPP
PTOPP is compatible with sampling-based optimal planners and shares common subroutines with sampling-based algorithms, such as the nearest neighbor search and edge collision check. This part details other primitive procedures PTOPP relies on, in particular, node attributes and primitive routines related to passage traversal status computation.
@@ 5.1 Path node attributes
Sampling-based planners maintain a graph or tree structure to organize samples. Without loss of generality, a sample $\mathbf{x}$ is referred to as a path/graph node whose position is drawn from a random sampling process. In OPP, basic node attributes include the predecessor (parent) and descendants/neighbors (children) in edge connections, the optimal path cost (cost) from the start node $\mathbf{x}_{0}$ to $\mathbf{x}$, and the path length (len). Attributes of specific data structures, e.g., pointers to the left and right child in kd-tree, are also assigned. In PTOPP, nodes additionally carry attributes of passage traversal and cell positioning status. passedWidths is the array of passage widths in the order traversed by the edge $(\mathbf{x}_{parent},\mathbf{x})$. The number of passages an edge passes through is generally small but may exceed one (see Figure 9). passedWidths is empty if no passage is crossed. sortedWidths corresponds to $\mathbf{p}_{\sigma}$ with $\sigma$ being the optimal path from $\mathbf{x}_{0}$ to $\mathbf{x}$ and is dynamically maintained as a linked list. cellIndex stores the Gabriel cell index where the node is located. In initialization, len, cost, and sortedWidths entries are assigned large values by default. Data Structure outlines these core path node attributes in C++ style.
@@ 5.2.1 Passage traversal check
Given a passage $\mathcal{P}_{i-j}$ and an edge $(\mathbf{x}_{1},\mathbf{x}_{2})$, the function Traverse$(\mathcal{P}_{i-j},\mathbf{x}_{1},\mathbf{x}_{2})$ returns True if $(\mathbf{x}_{1},\mathbf{x}_{2})$ passes $\mathcal{P}_{i-j}$ and False otherwise. While $\mathcal{P}_{i-j}$ represents a region, this examination is carried out between $l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})$ and $(\mathbf{x}_{1},\mathbf{x}_{2})$ so that only binary results exist. In 3D space, $l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})$ with its height interval becomes a plane similar to a virtual gate. Under the general position assumption of nodes, corner cases of colinearity or coplanarity between passage segments and edges are not considered. Passage traversal check is essentially equivalent to an intersection check as follows
@@
1 Struct Node:
@@ 3 Node *parent, *leftChild, *rightChild;
      4 list$<$Node*$>$ children (or adjacencyList);
@@
      7 array$<$double$>$ passedWidths;
      8 list$<$double$>$ sortedWidths;
@@
1 Input $\mathbf{x}_{1},\mathbf{x}_{2}$ with known $\mathbf{x}_{1}.$cellIndex and $l(\mathbf{x}_{1},\mathbf{x}_{2})$ being collison-free;
@@ 3 $C\leftarrow\emptyset;P\leftarrow\emptyset$;
            16 $\mathcal{P}\leftarrow$ the newly passed passage in $P_{t}\backslash P$;
@@ 18 $P\leftarrow P\cup P_{t};\,C\leftarrow C\cup\{c_{id}\}$;
Algorithm 6 PositionNode$(\mathbf{x}_{1},\mathbf{x}_{2})$
5.2.2 Passage traversal check by node positioning in Gabriel cells
For a collision-free edge $(\mathbf{x}_{1},\mathbf{x}_{2})\subset\mathcal{X}_{free}$ and the Gabriel cell $\mathcal{C}_{1}$ in which $\mathbf{x}_{1}$ is located ($\mathbf{x}_{1}\in\mathcal{C}_{1}$), the function SameCell$(\mathbf{x}_{1},\mathbf{x}_{2})$ returns True if $\mathbf{x}_{2}$ also lies inside $\mathcal{C}_{1}$ ($\mathbf{x}_{2}\in\mathcal{C}_{1}$) and False otherwise. This procedure fast examines if $(\mathbf{x}_{1},\mathbf{x}_{2})$ passes through any passage because only passages enclosing $\mathcal{C}_{1}$ need to be tested. Namely, the following holds
where $\mathcal{E}_{1}$ is the set of passage edges enclosing $\mathcal{C}_{1}$. Clearly, arguments are commutative, i.e., SameCell$(\mathbf{x}_{1},\mathbf{x}_{2})$ = SameCell$(\mathbf{x}_{2},\mathbf{x}_{1})$. $\mathcal{C}_{1}$ usually contains edges from obstacle boundaries that may collide with $(\mathbf{x}_{1},\mathbf{x}_{2})$. (28) holds due to the premise that $(\mathbf{x}_{1},\mathbf{x}_{2})$ is collision-free, which excludes the possibility that $(\mathbf{x}_{1},\mathbf{x}_{2})$ collides with an obstacle, passes no passages, and $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ belongs to two different cells. The illustration of various placements for an edge in cells is shown in Figure 9.
@@ 5.2.2 Passage traversal check by node positioning in Gabriel cells
Based on this routine, node positioning in Gabriel cells is completed rapidly. Nodes need to be located in cells, i.e., determining the cellIndex attribute, to retrieve passages enclosing them. Without iteratively examining all cells or relying on complicated point location query algorithms such as trapezoidal maps (Mark et al., 2008), the proximity of adjacent path nodes enables fast sequential node positioning. Suppose $\mathbf{x}_{1}$ is previously located and a new node $\mathbf{x}_{2}$ is linked to $\mathbf{x}_{1}$ through edge $(\mathbf{x}_{1},\mathbf{x}_{2})$. To determine $\mathbf{x}_{2}.$cellIndex, it suffices to investigate what cells $(\mathbf{x}_{1},\mathbf{x}_{2})$ traverses and the cells to be tested are thus limited. As in Algorithm 6, starting from the cell $\mathbf{x}_{1}$ is inside, the script examines each Gabriel cell in the order $(\mathbf{x}_{1},\mathbf{x}_{2})$ traverses until reaching the one $\mathbf{x}_{2}$ is located in. Since the edge is generally short compared with cell sizes, only a small number of cells will be passed before termination, resulting in an efficient node positioning and passage traversal check procedure. Meanwhile, other related attributes of $\mathbf{x}_{2}$ are also determined in this positioning procedure as in Algorithm 6.
5.2.3 Cost computation
Denote the optimal path from the start $\mathbf{x}_{0}$ to $\mathbf{x}$ online as $\sigma^{\prime}_{\mathbf{x}}$. Cost($\mathbf{x}$) returns the path cost from $\mathbf{x}$ attributes, i.e., $\texttt{Cost}(\mathbf{x})=f_{c}(\sigma^{\prime}_{\mathbf{x}})$. This cost is stored in $\mathbf{x}.\texttt{cost}$. $l(\mathbf{x}_{1},\mathbf{x}_{2})$ represents the straight-line path from $\mathbf{x}_{1}$ to $\mathbf{x}_{2}$. Suppose $l(\mathbf{x}_{1},\mathbf{x}_{2})$ is collision-free, NewCost$(\mathbf{x}_{1},\mathbf{x}_{2})$ computes the cost of the concatenated path $\sigma^{\prime}_{\mathbf{x}_{1}}|l(\mathbf{x}_{1},\mathbf{x}_{2})$ as the cost to come of $\mathbf{x}_{2}$ through $\mathbf{x}_{1}$. As $l(\mathbf{x}_{1},\mathbf{x}_{2})$ is newly concatenated, the variables used in cost computation are first updated. For example, a temporary sorted passage width vector is assigned $\mathbf{p}^{\prime}=\mathbf{p}_{\sigma^{\prime}_{\mathbf{x}_{1}}}$ for the joined path. $\mathbf{p}^{\prime}$ needs to be updated if $l(\mathbf{x}_{1},\mathbf{x}_{2})$ has traversed some passages. Then the cost is calculated according to the specific PTOPP type. See the detailed procedure in Algorithm 7 where the cost type is easily configurable.
@@ 5.2.4 Subtree node attributes update
A key operation to attain path optimality in many sampling-based planners is modifying edge connections locally in the neighborhood of newly sampled nodes, e.g., the rewire routine in RRT and its variants. The effect of changing a node’s edge connection propagates to its descendant nodes. Assume the edge $(\mathbf{x}_{1},\mathbf{x}_{2})$ is newly established by changing $\mathbf{x}_{2}$’s parent to $\mathbf{x}_{1}$. This modification influences $\mathbf{x}_{2}$’s descendants since optimal paths to them change accordingly. $\mathbf{x}_{2}$ is the root of the subtree composed of its descendant nodes. Attributes related to the optimal path to subtree nodes are updated using the new path $\sigma^{\prime}_{\mathbf{x}_{1}}|l(\mathbf{x}_{1},\mathbf{x}_{2})|\sigma^{\prime}(\mathbf{x}_{2}\sim\mathbf{x})$, where $\sigma^{\prime}(\mathbf{x}_{2}\sim\mathbf{x})$ represents the path segment from $\mathbf{x}_{2}$ to $\mathbf{x}$ in the subtree. This can be done by traversing the subtree with breadth-first search (BFS) or DFS. SubtreeUpdate() in Algorithm 8 details this procedure in a BFS backbone to update subtree node attributes of len, cost, and sortedWidths.
1 Input $\mathbf{x}_{1},\mathbf{x}_{2}$ with $l(\mathbf{x}_{1},\mathbf{x}_{2})$ being collison-free;
@@ 6 Insert passage widths in $P_{t}$ to $\mathbf{p}^{\prime}$;
Algorithm 7 NewCost$(\mathbf{x}_{1},\mathbf{x}_{2})$
1 Input $\mathbf{x}_{1},\mathbf{x}_{2}$ where $\mathbf{x}_{1}$ is the new parent of $\mathbf{x}_{2}$;
@@ 5 PositionNode$(\mathbf{x}_{1},\mathbf{x}_{2})$;
Algorithm 8 SubtreeUpdate$(\mathbf{x}_{1},\mathbf{x}_{2})$
6 Sampling-based algorithms for PTOPP
This section elaborates on sampling-based optimal planners for PTOPP integrating previous components. In particular, algorithmic integration with the two arguably most influential asymptotically optimal planners, optimal PRM (PRM) and optimal RRT (RRT) (Orthey et al., 2023), is detailed. The integration principles with their major planner variants are also briefly discussed.
6.1 PRM and variants for PTOPP
Primarily aimed at multi-query applications, PRM consists of the preprocessing phase constructing a roadmap from random samples in $\mathcal{X}_{free}$ and the query phase searching for the optimal path in this roadmap. It is probabilistically complete in finding feasible but non-optimal paths. To be asymptotically optimal, PRM enhances roadmap construction by adopting a connection radius as a function of the total sample number. In PTOPP, roadmap construction remains unchanged to capture the connectivity of $\mathcal{X}_{free}$. The key to achieving the optimal path lies in modifying the path search procedure. Let the undirected graph $G=(V,E)$ with vertex set $V$ and edge set $E$ represent the roadmap. Before path searching, $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ are first linked the $G$. The nodes $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ connect to are found by minimizing the specified cost $f_{c}(\cdot)$ under path length regulation. Suppose $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ are linked to $\mathbf{x}_{G,0},\mathbf{x}_{G,g}\in V$ respectively, then
$\begin{split}&\mathbf{x}_{G,0}=\operatorname*{arg\,min}_{\mathbf{x}\in V}\,\text{Len}(l(\mathbf{x}_{0},\mathbf{x}))\\ \operatorname{s.t.}\;\;&f_{c}(l(\mathbf{x}_{0},\mathbf{x}_{G,0}))=\min f_{c}(l(\mathbf{x}_{0},\mathbf{x}))\\ &\qquad l(\mathbf{x}_{0},\mathbf{x}_{G,0})\subset\mathcal{X}_{free}.\\ \end{split}$ (29)
$\mathbf{x}_{G,g}$ is obtained analogously by replacing $l(\mathbf{x}_{0},\mathbf{x}_{G,0})$ with $l(\mathbf{x}_{G,g},\mathbf{x}_{g})$. Alternatively, $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ may also be directly added to the roadmap as nodes. The query phase searches for the optimal path from $\mathbf{x}_{G,0}$ to $\mathbf{x}_{G,g}$ in $G$. Because of the path cost monotonicity in PTOPP, $G$ is a non-negatively weighted graph. Therefore, optimal path finding equals the shortest path problem where edge weights depend on passages passed by edges. Shortest path finding techniques in weighted graphs can be adopted readily with modifications to compute node costs in PTOPP.
The PTOPP pipeline built on PRM is outlined in Algorithm 9. Passage and cell preprocessing is carried out as in Section 3. $r_{PRM^{*}}$ represents the connection radius as a function of the batch sample number $n$ (Karaman and Frazzoli, 2011), i.e., $r_{PRM^{*}}=\gamma_{PRM}(\log(n)/n)^{1/d}$ where $\gamma_{PRM}>\gamma_{PRM}^{*}=2(1+1/d)^{1/d}(\mu(\mathcal{X}_{free})/\xi_{d}))^{1/d}$. $d=\text{dim}(\mathcal{X})$ is the dimension of the space $\mathcal{X}$. $\mu(\mathcal{X}_{free})$ and $\xi_{d}$ denotes the volume of $\mathcal{X}_{free}$ and the unit ball in the $\mathbb{R}^{d}$ Euclidean space, respectively. Separate node localization in Gabriel cells is only performed for $\mathbf{x}_{G,0}$ initially. The query phase adopts Dijkstra’s algorithm with two core substitute subroutines. First, the temporary path cost through node $\mathbf{u}$ to its neighbor $\mathbf{v}$ is calculated by NewCost$(\mathbf{u,v})$. Second, if the temporary path cost is smaller than $\mathbf{v}.$cost, SubtreeUpdate$(\mathbf{u,v})$ is invoked to update path connections and node attributes. In PRM, nodes maintain adjacent nodes, not children, SubtreeUpdate$(\mathbf{u,v})$ thus terminates without entering the BFS procedure. Note that under path length regulation, node precedence is compared lexicographically by the tuple (cost, len). $\mathbf{u}$ is considered to have a smaller PTOPP path cost than $\mathbf{v}$ if
@@ 6.1 PRM and variants for PTOPP
This is utilized when determining the minimum cost node (line 9 in Algorithm 9 with a priority queue) and comparing the temporary cost with node cost (line 9 in Algorithm 9). At last, path nodes are retrieved by recursively fetching nodes’ predecessor parent, reversing its order in GetPath(), and concatenating $\mathbf{x}_{0},\mathbf{x}_{g}$.
The underlying architecture of PRM for PTOPP depicts a generic pipeline of graph-based solution comprising roadmap construction and optimal path search. It applies to other optimal PRM variants, such as simplified PRM (sPRM), that only differ in the roadmap construction phase. In addition, one important branch is pure cell-based methods enabled by cell decomposition (Choset et al., 2005; LaValle, 2006). They are particularly promising in PTOPP because no extra cell decomposition step is required. Rather than random samples, the roadmap uses sparse non-random samples in cells, e.g., centroids and cell edge midpoints, as vertices to capture the connectivity of divided regions. Because of the dramatic decrease in node number, pure cell-based methods enjoy high computational efficiency. The inherent downside is that the found path is coarse and subject to resolution optimality. Heavy path postprocessing is usually needed. It is also complicated to extend to high-dimensional spaces.
1 Input $\mathbf{x}_{0},\mathbf{x}_{g}$ and conduct passage and cell preprocessing;
@@
2 $V\leftarrow\{\texttt{SampleFree}()\}_{i=1,...,n};E\leftarrow\emptyset$;
3 foreach $\mathbf{v}\in V$ do
      4 $V_{near}\leftarrow\texttt{Near}(G=(V,E),\mathbf{v},r_{PRM^{*}})\backslash\{\mathbf{v}\}$;
      5 foreach $\mathbf{u}\in V_{near}$ do
            6 if CollisionFree$(\mathbf{v,u})$ then
                  7 $E\leftarrow E\cup\{(\mathbf{v,u}),(\mathbf{u,v})\}$;
@@
[…]

### v2

@@ 4.3 Constrained passage width PTOPP
The last category of PTOPP focuses on problems with constraints on passed passage widths, termed constrained passage width PTOPP (CPW-PTOPP). The most representative case is that there exists a minimum admissible passage width $\underline{\varepsilon}_{p}$ to accommodate robots, similar to the preset clearance requirement. Technically, this is solvable by labeling passage regions of widths smaller than $\varepsilon_{p}$ as obstacles to prohibit traversing. In PTOPP, such constraints are readily imposed on path costs without the labeling processing above. Firstly, CPW-PTOPP can be easily integrated into shortest path planning. If the path passes a passage narrower than $\underline{\varepsilon}_{p}$, a penalty is imposed by adding a large value $K_{p}$ to the path cost, which becomes
@@
where $|\mathbf{p}_{\sigma}|_{\underline{\varepsilon}_{p}}$ is the number of widths in $\mathbf{p}_{\sigma}$ smaller than $\underline{\varepsilon}_{p}$. The cost compatibility is easy to verify.
Next, constrained passage width could be enforced in PTOPP categories by truncating passage widths smaller than $\underline{\varepsilon}_{p}$ to be a negative value (e.g., $1/\|l(\mathbf{p}_{i}^{*},\mathbf{p}_{j}^{*})\|$ here) so that widths in $\mathbf{p}_{\sigma}$ are updated as follows
@@
This causes the path cost $f_{c}(\sigma)$ to soar after passing passages narrower than $\underline{\varepsilon}_{p}$, excluding them from the optimal path’s route. $f_{c}(\sigma)$ refers to any cost above. Hence, CPW-PTOPP is adapted from previous PTOPP categories by simply applying (26), making passage width constraints quite convenient to enforce. Note that both MPW-PTOPP and GPW-PTOPP maximize the smallest passage width $f_{p}(\sigma)$ under the proposed costs. If $f_{p}(\sigma)>\underline{\varepsilon}_{p}$, (26) does not change the optimal path. If $f_{p}(\sigma)\leq\underline{\varepsilon}_{p}$, (26) helps select wider passages and decrease the number of passages narrower than $\underline{\varepsilon}_{p}$ for a smaller cost.
@@ 5 Primitive procedures in PTOPP
PTOPP shares common subroutines with sampling-based optimal planners such as the nearest neighbor search and collision check. This part details other primitive procedures PTOPP relies on, in particular, node attributes and primitive routines for passage traversal status computation.
@@ 5.1 Path node attributes
Sampling-based planners maintain a graph or tree to organize samples. A sample $\mathbf{x}$ is referred to as a path/graph node whose position is drawn from a random sampling process. In OPP, basic node attributes include the predecessor (parent) and descendants/neighbors (children), the optimal path cost (cost) from the start node $\mathbf{x}_{0}$ to $\mathbf{x}$, and the associated path length (len). Attributes of other data structures, e.g., left/right child in kd-tree, are also assigned. In PTOPP, nodes additionally carry attributes of passage traversal and cell positioning status. passedWidths is the array of passage widths in the order traversed by the edge $(\mathbf{x}_{parent},\mathbf{x})$. The number of passages an edge passes through is generally small but may exceed one (see Figure 9). sortedWidths corresponds to $\mathbf{p}_{\sigma}$ with $\sigma$ being the optimal path from $\mathbf{x}_{0}$ to $\mathbf{x}$ and is dynamically maintained as a linked list. cellIndex stores the Gabriel cell index where the node is located. In initialization, sortedWidths entries are assigned large values. Data Structure outlines these core path node attributes in C++ style.
@@ 5.2.1 Passage traversal check
[…]
@@
5.2.2 Passage traversal check by node positioning in Gabriel cells
[…]
where $\mathcal{E}_{1}$ is the set of passage edges enclosing $\mathcal{C}_{1}$. $\mathcal{C}_{1}$ usually contains edges from obstacle boundaries. (28) holds due to the premise that $(\mathbf{x}_{1},\mathbf{x}_{2})$ is collision-free, which excludes the possibility that $(\mathbf{x}_{1},\mathbf{x}_{2})$ collides with an obstacle, passes no passages, and $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ belongs to two different cells. The illustration of various placements for an edge in cells is shown in Figure 9.
Based on this routine, node positioning in Gabriel cells is completed rapidly. Nodes need to be located in cells, i.e., determining the cellIndex attribute, to retrieve passages enclosing them. Without iteratively examining all cells or relying on complicated point location query algorithms such as trapezoidal maps (mark2008computational), the proximity of adjacent path nodes enables fast sequential node positioning. Suppose $\mathbf{x}_{1}$ is previously located and a new node $\mathbf{x}_{2}$ is linked to $\mathbf{x}_{1}$ through edge $(\mathbf{x}_{1},\mathbf{x}_{2})$. To determine $\mathbf{x}_{2}.$cellIndex, it suffices to investigate what cells $(\mathbf{x}_{1},\mathbf{x}_{2})$ traverses and the cells to be tested are thus limited. As in Algorithm 6, starting from the cell $\mathbf{x}_{1}$ is inside, the script examines each Gabriel cell in the order $(\mathbf{x}_{1},\mathbf{x}_{2})$ traverses until reaching the one $\mathbf{x}_{2}$ is located in. Since the edge is generally short compared with cell sizes, only a small number of cells will be passed before termination, resulting in an efficient node positioning and passage traversal check procedure. Meanwhile, other related attributes of $\mathbf{x}_{2}$ are determined in this positioning procedure as in Algorithm 6.
1Struct Node:
@@ 3 Node *parent, *leftChild, *rightChild;
 4 list<<Node*>> children (or adjacencyList);
@@
 7 array<<double>> passedWidths;
 8 list<<double>> sortedWidths;
@@
1Input $\mathbf{x}_{1},\mathbf{x}_{2}$ with known $\mathbf{x}_{1}.$cellIndex and $l(\mathbf{x}_{1},\mathbf{x}_{2})$ being collison-free;
@@ 3 $C\leftarrow\emptyset;P\leftarrow\emptyset$;
    16 $\mathcal{P}\leftarrow$ the newly passed passage in $\PP_{t}\backslash P$;
@@ 18 $P\leftarrow P\cup P_{t};\,C\leftarrow C\cup\{c_{id}\}$;
[…]
5.2.3 Cost computation
[…]
@@ 5.2.4 Subtree node attributes update
A key operation to attain path optimality in many sampling-based planners is modifying edge connections locally in the neighborhood of new samples, e.g., the rewire routine in RRT. The effect of changing a node’s edge connection propagates to its descendants. Assume edge $(\mathbf{x}_{1},\mathbf{x}_{2})$ is newly established by changing $\mathbf{x}_{2}$’s parent to $\mathbf{x}_{1}$. Paths to $\mathbf{x}_{2}$’s descendants change accordingly. $\mathbf{x}_{2}$ is the root of the subtree composed of its descendant nodes. Subtree nodes’ path attributes are updated using the new path $\sigma^{\prime}_{\mathbf{x}_{1}}|l(\mathbf{x}_{1},\mathbf{x}_{2})|\sigma^{\prime}(\mathbf{x}_{2}\sim\mathbf{x})$, where $\sigma^{\prime}(\mathbf{x}_{2}\sim\mathbf{x})$ represents the path from $\mathbf{x}_{2}$ to $\mathbf{x}$ in the subtree. This is done by traversing the subtree with breadth-first search (BFS) or DFS. SubtreeUpdate() in Algorithm 8 shows a BFS backbone to update subtree node attributes of len, cost, and sortedWidths.
6 Sampling-based algorithms for PTOPP
This part elaborates on sampling-based optimal planners for PTOPP integrating previous components. Integration with the two arguably most influential asymptotically optimal planners, optimal PRM (PRM) and optimal RRT (RRT), is detailed. Integration principles with their major variants are also briefly discussed.
6.1 PRM and variants for PTOPP
Primarily aimed at multi-query applications, PRM consists of the preprocessing phase constructing a roadmap from random samples in $\mathcal{X}_{free}$ and the query phase searching for the optimal path in this roadmap. In PTOPP, roadmap construction remains unchanged to capture the connectivity of $\mathcal{X}_{free}$. The key to achieving the optimal path lies in modifying the path search procedure. Let the undirected graph $G=(V,E)$ with vertex set $V$ and edge set $E$ represent the roadmap. The nodes in $G$ that $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ connect to are first found by minimizing the cost $f_{c}(\cdot)$ under path length regulation. Suppose $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ are linked to $\mathbf{x}_{G,0},\mathbf{x}_{G,g}\in V$ respectively. Alternatively, $\mathbf{x}_{0}$ and $\mathbf{x}_{g}$ may also be directly added to the roadmap as nodes. The query phase searches for the optimal path from $\mathbf{x}_{G,0}$ to $\mathbf{x}_{G,g}$ in $G$. Because of the path cost monotonicity in PTOPP, $G$ is a non-negatively weighted graph. Therefore, optimal path finding equals the shortest path problem where edge weights depend on passages passed by edges. Shortest path finding techniques in weighted graphs can be adopted readily with modifications to compute node costs in PTOPP.
1Input $\mathbf{x}_{1},\mathbf{x}_{2}$ with $l(\mathbf{x}_{1},\mathbf{x}_{2})$ being collison-free;
@@ 6 Insert passage widths in $P_{t}$ to $\mathbf{p}^{\prime}$;
[…]
1Input $\mathbf{x}_{1},\mathbf{x}_{2}$ where $\mathbf{x}_{1}$ is the new parent of $\mathbf{x}_{2}$;
@@
[…]
@@
This is utilized when determining the minimum cost node (line 9 in Algorithm 9) and comparing the temporary cost with node cost (line 9 in Algorithm 9).
PRM for PTOPP depicts a generic graph-based solution framework comprising roadmap construction and optimal path search. It applies to other optimal PRM variants, e.g., simplified PRM (sPRM), which only differ in their roadmap construction. In addition, one important branch is pure cell-based methods enabled by cell decomposition (choset2005principles; lavalle2006planning). They are particularly promising in PTOPP because no extra cell decomposition step is required. Rather than random samples, the roadmap uses sparse non-random samples in cells, e.g., centroids and edge midpoints, as vertices to capture the connectivity of divided regions. Thanks to the dramatic decrease in node number, pure cell-based methods enjoy high computational efficiency, though heavy path postprocessing is usually needed.
@@
1 Input $\mathbf{x}_{0},\mathbf{x}_{g}$ and conduct passage and cell preprocessing.;
2 Construct $G=(V,E)$ via random sampling.;
@@
3 $\mathbf{x}_{G,0},\mathbf{x}_{G,g}\leftarrow X;Q\leftarrow V;\mathbf{x}_{G,0}.\texttt{cost}\leftarrow\texttt{Cost}(\mathbf{x}_{G,0})$;